# Preparations for Quantum Computing in Hadron Physics

### v1

@@ 7 Progress towards the Equation of State of Neutron Stars
in the low temperature limit this reduces to a ground-state-search problem, for which algorithms such as VQE could be used. The tricky part is then to construct suitable ansätze, which will necessary depend on the concrete realization of QCD; a blind variational search will be probably plagued with barren plateaus.
Our group is exploring an approach to calculate the ground-state energy from the particle encoding in section 5 within Weyl-gauged QCD, with a simple Hamiltonian but the need to enforced Gauss’ law independently, see subsection 6.1. We count on making a proof of principle available soon.
@@ 8 Progress towards Fragmentation Functions
For this, we have completely recalculated the Hamiltonian of Quantum Chromodynamics in Light Front Quantization and in Light Front Gauge [126]. We adopted a jet initiated by a charm quark with a large starting longitudinal light-front momentum $p_{j}^{+}=\Lambda$ that sets the maximum of our longitudinal grid, so that the emitted $J/\psi$ carries a fraction $z$ given as $p_{h}^{+}=zp_{j}^{+}$. We employed three qubits (eight longitudinal momentum states) per particle, and neglected all transverse momenta (in principle not a terrible approximation in jet fragmentation where $p_{\perp}\ll p_{+}$). To this we need to add color, flavor if the parton is a quark or antiquark, and spin, to total 7–8 qubits per parton.
@@
A few calculational details are in order. First, to extract the Fragmentation Function from Eq. (2) we make an ansatz $J/\psi$ longitudinal distribution,
$|J/\Psi\rangle=\sum\delta_{c_{q}c_{\bar{q}}}\frac{\chi_{0}(x)\,\vec{\sigma}_{ij}}{\sqrt{x\,(z-x)}}\left|x\ i\,c_{q},(z-x)\ j\,c_{\bar{q}}\right\rangle\ \ \$(74)
@@
    Given enough memory, the number of particles would become large (“saturation regime”) and entropy would be maximized once the probability would be spread among the many degrees of freedom. In our simulation, with at most four particles, the unitary operator can start cycling back. We compute $S$ and look for a plateau.
@@
The outcome of an example computation is shown in Fig. 7. The grid cannot be very tight (due to the small number of qubits which can be simulated), the maximum number of particles is 4 (up to two quarks, an antiquark and a gluon) and either the group was shortened to $SU(2)$ or not all terms of the Hamiltonian (particularly seagulls and fork interactions germane to the light-front quantization) were employed in all calculations; but in spite of these limitations, necessary to obtain finite running times on a standard computer cluster, we believe that we have demonstrated the feasibility of extracting fragmentation functions from upcoming quantum computers with a beginning-to-end algorithm that serves as a demonstrator.
Figure 7: Fragmentation function $D_{c}^{J/\psi}$ extracted from a classical simulation of a quantum computer with order 30 qubits. An $SU(3)$ calculation (stars) was run with only part of the interaction vertices (splitting, absorbing and pair creation) and one can see, from the equivalent $SU(2)$ computation at equal $\alpha_{s}$ and conditions (squares) the dependence on the color factor. The entire Hamiltonian was run only in the $SU(2)$ case (diamonds) but the difference is seen to be small. These three computations had a meager $N=4$ partition of the longitudinal momentum; an additional $N=8$ computation (circles) is shown to indicate the sensitivity thereto.
@@ 9 Progress towards time evolution in hadron physics
The first one [76] is an $SU(3)$ gauge theory with the canonically quantized lattice Hamiltonian already described (section 4). An interesting method employed to mitigate quantum computer error consisted in computing a physics run evolving forward for $N_{T}$ Trotter time steps; and then calculate again, evolving forward for $N_{T}/2$ steps and backwards for another $N_{T}/2$, with the aim of reconstructing the initial state, whose comparison with the original preparation allows to characterize the computer error. The number of time steps $N_{T}$ was of order 4 to 8 depending on the machine employed. The evolution of both the quark + antiquark particle number (shown in the figure) as well as that of the gauge field were tracked.
@@
Another example of this type of calculation, again for QCD in 1+1 dimensions, can be found in [77], the Kogut-Sussking Hamiltonian of Eq. 16 is formulated in the axial gauge $A^{x}=0$, allowing for the gauge degrees of freedom to be transformed into long-range interactions between quarks. This generates “color edge” states at the end of the lattice which are artifically “pulled up” in energy adding to the Hamiltonian a term that depends on the sum of color charges. The problem is formulated on a lattice with quarks and antiquarks of two flavours, requiring $3\times 2\times 2=12$ qubits per lattice site. The low-energy spectrum is solved classically for several values of the coupling constant $g$ and with a quantum annealer, a special device to solve variational problems. Later the authors use the Jordan-Wigner transform to write the Hamiltonian operator in terms of Pauli operators, and develop efficient circuits for the simulation of real-time evolution. The results, which closely resemble exact evolution, are obtained with aid of error-mitigation techniques. Closely related is[78], in which the authors add to the Hamiltonian an effective four lepton interaction to generate baryon $\beta$ decays. The simulation is run on a $20$ trapped-ion quantum computer[145] filtering the results that are obvious errors (post-selection) without error mitigation, obtaining results at percente level.
@@ 10.1 Fragmentation in the NJL model and the Schwinger model
$\mathcal{L}=\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m)\psi+g(\bar{\psi}\psi)^{2}$ (76)
@@
$e^{-i\eta K}e^{-iHt}\ \overline{\psi}(0,\pm x^{3})e^{iHt}e^{i\eta K}=\overline{\psi}(-\gamma(t\pm vx^{3}),\gamma(vt\pm x^{3})e^{-\frac{1}{2}\eta\gamma^{5}})$ (78)
@@ 10.2 Jet energy loss and entropy production
Figure 11: Heat–map type plots of the entropy produced by a jet (coming in from the top left in each plot) scattering in a slice of a medium (center of the figure) for different values of the total charge of the jet in the Schwinger model of Eq. (79) Reprinted[147] under the terms of the Creative Commons 4.0 License, https://creativecommons.org/licenses/by/4.0/
@@ 10.3 State preparation and energy loss in charged medium
Research on state preparation concerns the construction of appropriate ansätze and the minimization of objective functions such as the expectation value of the energy or the infidelity with respect to a desired target state. Theoretical and practical inputs about the process are based on the preparation and evolution within simplified models, such as the the 1+1 D Schwinger model[79], in which the “hadrons” are $e^{+}-e^{-}$ bound states, exemplified with the use of the $112$ qubits quantum-computer ibm_torino[95]. The vacuum of the theory is prepared first, and subsequently a “hadron” on top of it, keeping track of the evolution of the chiral condensate expectation value as the system evolves in time for up to $14$ Trotter steps. The state-preparations procedure is named SC-ADAPT-VQE and consists on repeatedly tuning selected unitary transformation with classical circuits for increasing lattice lengths until convergence, finally extrapolating this transformations to the real length on the quantum device. It is found that classical simulators can manage the preparation for quite large lengths as the correlations between physical sites remain small, but run intro troubles for time evolution.
@@
The preparation and time-evolution circuits of[79] are combined with the presence of static charges to mimic, via classical simulators, the evolution of heavy particles in dense media [80]. Lattice effects are probed through the evolution of an initial bound state, where the light dynamical degrees of freedom couple to a heavy charge. Due to the lattice dispersion relations, once a critical velocity $v^{*}$, less than the speed of light, is exceeded, the light degrees of freedom decouple from the charge and excite the lattice in a process resembling hadronization. To suppress such effects, the velocity of the moving charge is kept small as it traverses the lattice populated with other static charges. Within this setup, the energy loss and entanglement are computed, allowing lattice discretization effects to be isolated. Entanglement is found to be more sensitive to lattice artifacts than other classical observables. In addition, quantum circuits and associated computational costs are discussed.
@@ 11 Outlook
There are dedicated reviews on Quantum Computing which cover topics in High-Energy Physics [72] and Nuclear Physics [149], but we feel that there is a need to point out the important niche applications in Hadron Physics, straddling both fields, sharing from the large scales of the first but also from the nonperturbative problems of the second. We have striven to give, in a few brushstrokes, a colorful picture of some interesting problems which can be addressed when quantum computers become a more practical tool, without precluding that several other interesting applications may be devised in the field.
@@
This entails that several extant works fall back onto topical phenomenological models addressing one or another aspect of the full theory. As hardware advances, the advantage of directly working with discretizations and truncations of QCD lies in the possibility of controlling systematic uncertainties through well-defined computational parameters—a feature that also applies to quantum simulations.
Finally, one crucial aspect that quantum computing practitioners entering the field of quantum field theory simulations should bear in mind is the renormalization, or flow, of the Hamiltonian parameters when the scale of the problem changes (and hence the discretization), whether in momentum space or particle-number space. The following paragraph is devoted to this issue.
Renormalization Group evolution
We have not dedicated space in this brief appraisal to the renormalization of the QCD Hamiltonian, but a comment is in order here, since, unlike the quantum mechanical theories often transcribed for a quantum computer, Chromodynamics will require serious work to handle its renormalization intricacies. A promising line of research is that of the Renormalization Group Procedure for Effective Particles (RGPEP) [155] recently deployed for a Yukawa Hamiltonian [136]. The traditional idea of continuously transforming the Hamiltonian via a spectrum-conserving unitary rotation parametrized by a control variable $s$ (for example, a scale),
$H(s)=U(s)H(0)U^{\dagger}(s)$ (80)
@@
New counterterms are then added to the Hamiltonian and the eigenvalues are extracted. This, or other RGE techniques [150, 151, 152] will need to be automated into the workflow for QCD on quantum computers, which is currently best described as simply cut-off (regulated).
The RGPEP is most often formulated using the Light Front, and the Hamiltonian requires renormalization. The infinite number of degrees of freedom requires truncation; to study the sensitivity thereto, calls for a renormalization-group method to be applied. The RGPEP enables the construction of effective Hamiltonians through a similarity transformation that depends on a scale parameter. The eigenvector of the effective Hamiltonian contains a small number of non-negligible Fock components, thereby reducing the complexity of the description. Such an approach appears promising for reducing the number of degrees of freedom without truncating terms that encode essential dynamics. Simulations with effective Hamiltonians have been performed for a Yukawa theory within RGPEP, showing that the computational cost to block encode the renormalized Hamiltonian is comparable to block encoding the bare Hamiltonian [135, 136].

### v2

@@ 7 Progress towards the Equation of State of Neutron Stars
in the low temperature limit this reduces to a ground-state-search problem, for which algorithms such as VQE could be used. The tricky part is then to construct suitable ansätze, which will necessarily depend on the concrete realisation of QCD; a blind variational search will be probably plagued with barren plateaus.
Our group is exploring an approach to calculate the ground-state energy from the particle encoding in section 5 within Weyl-gauged QCD, with a simple Hamiltonian but the need to enforced Gauss’ law independently, see subsection 6.1. To be more specific, we divide the system into small and independent cells with periodic boundary conditions in which the final aim is to minimize the cost function
$E(\vec{\theta};\mu_{f},\nu)=\left\langle\psi(\vec{\theta})\right|\hat{H}+\mu_{f}\hat{N}_{f}+\nu\,\hat{\mathcal{G}}^{2}\left|\psi(\vec{\theta})\right\rangle\ ,$ (74)
on a quantum memory with a few dozens of particles. $\hat{H}$ is the QCD Hamiltonian in the Weyl-gauge, $\hat{\mathcal{G}}^{2}$ is the Gauss operator squared [^3] With $\mathcal{G}$ a self-adjoint operator, $\left\langle\psi\right|\mathcal{G}^{2}\left|\psi\right\rangle=\sum_{\phi}\left\langle\psi\right|\mathcal{G}\left|\phi\right\rangle\left\langle\phi\right|\mathcal{G}^{\dagger}\left|\psi\right\rangle=\sum_{\phi}|\left\langle\psi\right|\mathcal{G}\left|\phi\right\rangle|^{2}\geq 0$: its minimization corresponds to $\langle\phi|\mathcal{G}=0$ which ensures the gauge choice by enforcing Gauss’s law. and summed over positions and colours, $\hat{N}_{f}$ is the number of $f$-flavoured fermions, $\vec{\theta}$ are the variational parameters, $\mu_{f}$ are the chemical potentials and $\nu$ is a Lagrange multiplier. The EoS can then be constructed by analysing the change in energy density with the volume of these cells and employing standard statistical physics. To give an example on the use of the encoding of sec. 5 to measure expectation values, consider the splitting part of the three gluon vertex in Weyl gauge,
$H_{g21}=g\int\left[p,k,q\right]f^{abc}\left(p^{j}\delta^{i}_{l}+k^{l}\delta^{i}_{j}+q^{j}\delta^{i}_{l}\right)\left(i\,a^{\dagger}_{pia}a^{\dagger}_{kjb}a_{qlc}+h.c.\right)\tilde{\delta}\left(k+q-p\right)\ ,$ (75)
with a similar term for the Gauss operator squared
$\mathcal{G}_{g21}^{2}=2g\int\left[kpq\right]f^{abc}$ $\left(E_{k}E_{p}p^{i}\delta_{lj}+E_{q}E_{p}q^{l}\delta_{ji}\right$.
$\left.+E_{q}E_{p}p^{i}\delta_{lj}\right)\left(ia_{pia}^{\dagger}a_{kjb}^{\dagger}a_{qlc}+h.c.\right)\ \tilde{\delta}\left(p+k-q\right)\ ,$ (76)
with $\int\left[k\right]=\int\frac{d^{3}k}{\left(2\pi\right)^{3/2}\sqrt{2E_{k}}}$ and $\tilde{\delta}\left(p\right)=\left(2\pi\right)^{3}\delta^{3}\left(p\right)$. If the variational state is properly symmetrized and the memory is filled in order, the product of Eqs. (43,45) simplify to:
${\left\langle{\psi}\right|}a_{\rho}^{\dagger}a_{\eta}^{\dagger}a_{\omega}+h.c.{\left|{\psi}\right\rangle}={\left\langle{\psi}\right|}\left\{\sum_{j=1}^{N_{f}}j\sqrt{j+1}\ {\text{$\mathbb{P}^{({1})}_{{0}}$}}\otimes\right$. $\left({\text{$\mathfrak{C}_{{1}{0}}$}}\otimes{\text{$\mathfrak{s}^{\dagger}_{{\rho}}$}}\right)_{j+1}\otimes$
$\left.\left({\text{$\mathfrak{C}_{{1}{1}}$}}\otimes{\text{$\mathfrak{s}^{\dagger}_{{\eta}}$}}{\text{$\mathfrak{s}_{{\omega}}$}}\right)_{j}\otimes{\text{$\mathbb{P}^{({1})}_{{1}}$}}+h.c.\right\}{\left|{\psi}\right\rangle}\ ,$ (77)
where $\rho,\eta,\omega$ are compact indices for the gluon quantum numbers. This expression provides a LCU decomposition of the Fock operator product in terms of unitary matrices, since the projectors $\mathbb{P}$, controls $\mathfrak{C}$ and set $\mathfrak{s}^{\dagger}$ and scrap $\mathfrak{s}$ operators can all be written in terms of products of Pauli matrices.
@@ 8 Progress towards Fragmentation Functions
For this, we have completely recalculated the Hamiltonian of Quantum Chromodynamics in Light Front Quantisation and in Light Front Gauge [126]. We adopted a jet initiated by a charm quark with a large starting longitudinal light-front momentum $p_{j}^{+}=\Lambda$ that sets the maximum of our longitudinal grid, so that the emitted $J/\psi$ carries a fraction $z$ given as $p_{h}^{+}=zp_{j}^{+}$. We employed three qubits (eight longitudinal momentum states) per particle, and neglected all transverse momenta (in principle not a terrible approximation in jet fragmentation where $p_{\perp}\ll p_{+}$). To this we need to add colour, flavour if the parton is a quark or antiquark, and spin, to total 7–8 qubits per parton.
@@
A few calculational details are in order. First, to extract the Fragmentation Function from Eq. (2) we make an ansatz $J/\psi$ longitudinal distribution
$|J/\Psi\rangle=\sum\delta_{c_{q}c_{\bar{q}}}\frac{\chi_{0}(x)\,\vec{\sigma}_{ij}}{\sqrt{x\,(z-x)}}\left|x\ i\,c_{q},(z-x)\ j\,c_{\bar{q}}\right\rangle\ ,$ (78)
@@
    Given enough memory, the number of particles would become large (“saturation regime”) and entropy would be maximised once the probability would be spread among the many degrees of freedom. In our simulation, with at most four particles, the unitary operator can start cycling back. We compute $S$ and look for a plateau.
@@
The outcome of an example computation is shown in Fig. 7. The grid cannot be very tight (due to the small number of qubits which can be simulated), the maximum number of particles is 4 (up to two quarks, an antiquark and a gluon) and either the group was shortened to $SU(2)$ or not all terms of the Hamiltonian (particularly seagulls and fork interactions germane to the light-front quantisation) were employed in all calculations; but in spite of these limitations, necessary to obtain finite running times on a standard computer cluster, we believe that we have demonstrated the feasibility of extracting fragmentation functions from upcoming quantum computers with a beginning-to-end algorithm that serves as a demonstrator.
Figure 7: Fragmentation function $D_{c}^{J/\psi}$ extracted from a classical simulation of a quantum computer with order 30 qubits. An $SU(3)$ calculation (stars) was run with only part of the interaction vertices (splitting, absorbing and pair creation) and one can see, from the equivalent $SU(2)$ computation at equal $\alpha_{s}$ and conditions (squares) the dependence on the colour factor. The entire Hamiltonian was run only in the $SU(2)$ case (diamonds) but the difference is seen to be small. These three computations had a meager $N=4$ partition of the longitudinal momentum; an additional $N=8$ computation (circles) is shown to indicate the sensitivity thereto.
Parton Distribution Functions (pdfs),
the distributions parametrising the inverse problem –that of extracting a parton constituent from a hadron instead of the fragmentation of a hadron from a jet– have also been an object of recent scrutiny. For example, in the Nambu-Jona-Lasinio model in 1+1 dimension [144], the Schwinger model [145, 146] (see subsection 10.1 below for a brief discussion thereof).
Typically, the resulting distribution functions in this type of constituent-like models are valence-like, without the characteristic low-$x$ growth revealed by HERA-age experiments [147]. One would expect that the ability of the quantum computer to one day manage large numbers of particles would allow to see the low–$x$ growth of parton distributions driven by multigluon emission. For the time being, the regime that quantum computers can access is the same that lattice gauge theory probes via moments or the LaMEFT (Large-Momentum Effective Field Theory), so that we do not see it as an immediate priority.
@@ 9 Progress towards time evolution in hadron physics
The first one [76] is an $SU(3)$ gauge theory with the canonically quantised lattice Hamiltonian already described (section 4). An interesting method employed to mitigate quantum computer error consisted in computing a physics run evolving forward for $N_{T}$ Trotter time steps; and then calculate again, evolving forward for $N_{T}/2$ steps and backwards for another $N_{T}/2$, with the aim of reconstructing the initial state, whose comparison with the original preparation allows to characterise the computer error. The number of time steps $N_{T}$ was of order 4 to 8 depending on the machine employed. The evolution of both the quark + antiquark particle number (shown in the figure) as well as that of the gauge field were tracked.
@@
Another example of this type of calculation, again for QCD in 1+1 dimensions, can be found in [77], the Kogut-Sussking Hamiltonian of Eq. 16 is formulated in the axial gauge $A^{x}=0$, allowing for the gauge degrees of freedom to be transformed into long-range interactions between quarks. This generates “colour edge” states at the end of the lattice which are artificially “pulled up” in energy adding to the Hamiltonian a term that depends on the sum of colour charges. The problem is formulated on a lattice with quarks and antiquarks of two flavours, requiring $3\times 2\times 2=12$ qubits per lattice site. The low-energy spectrum is solved classically for several values of the coupling constant $g$ and with a quantum annealer, a special device to solve variational problems. Later the authors use the Jordan-Wigner transform to write the Hamiltonian operator in terms of Pauli operators, and develop efficient circuits for the simulation of real-time evolution. The results, which closely resemble exact evolution, are obtained with aid of error-mitigation techniques. Closely related is[78], in which the authors add to the Hamiltonian an effective four lepton interaction to generate baryon $\beta$ decays. The simulation is run on a $20$ trapped-ion quantum computer[148] filtering the results that are obvious errors (post-selection) without error mitigation, obtaining results at percent level.
@@
9.1 Jet broadening
We now briefly discuss a calculation [150] within the light-front Hamiltonian formalism showcasing three-dimensional jet evolution in a medium, which demonstrates transverse momentum broadening. The interest in having a quantum computer assess this problem is that perturbative calculations exist but are supposed to be valid at large temperatures and virtualities of the jet’s leading parton. However, many observations of jets do not satisfy those stringent conditions and perturbation theory is more questionable, with phenomenological approaches [151] suggesting larger values of the quenching parameter. Once more, while there are lattice gauge theory computations [153] attempting to extract $\hat{q}$, the jet quenching parameter, this is a quantity defined along a lightlike path, and in an Euclidean computation some form of analytical extrapolation on noisy data is needed, entailing large uncertainties. In fact, for temperatures below 300 MeV, the data there reported is not very meaningful. So this can be another project where quantum computers make headway.
The Light-front Hamiltonian for the parton probe driving the jet is split [150] as $P^{-}=\frac{{\bf p}_{\perp}^{2}}{2p^{+}}+gT^{a}A^{a\ -}$ (because other components of $A$ are suppressed by powers of $p^{+}$, the large scale). The dominant component is then taken to satisfy the classical Yang-Mills equations with a local source. The computed physical quantity, the quenching parameter $\hat{q}$, is defined from the broadening of the jet (per unit length $L$) in transverse momentum,
$\hat{q}=\frac{\langle{\bf p_{\perp}^{2}(L)}\rangle-\langle{\bf p_{\perp}^{2}(0)}\rangle}{L}\ .$ (80)
The algorithms employed are akin to those described above, so we refer the reader to the original publication [150]. Due to memory restrictions with a small number of qubits, the gauge groups chosen were $U(1)$ and $SU(2)$. For the latter, closer to QCD, the quenching parameter is obtained to be about $\hat{q}\simeq 0.24$GeV3 at a scale $Q_{s}^{2}\simeq 5$GeV2 and with a gluon masslike parameter $m_{g}\simeq 0.1$ GeV (which must give an indication of the scale of the temperature if the medium is thought of as thermal).
Jet energy loss is also discussed in a 1+1 Schwinger model with a wave packet plowing through a static medium in subsec.  (10.2) below. Because of the one-dimensional setup, particles can separate only in the longitudinal direction instead of opening a cone with transverse momentum, so many features of jet physics are absent. Other approaches to study jet propagation include the use of the open quantum system formalism with a scalar field-theory mock of the QGP medium [152].
@@ 10.1 Fragmentation in the NJL model and the Schwinger model
$\mathcal{L}=\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m)\psi+g(\bar{\psi}\psi)^{2}\ ,$ (81)
@@
$e^{-i\eta K}e^{-iHt}\ \overline{\psi}(0,\pm x^{3})e^{iHt}e^{i\eta K}=\overline{\psi}(-\gamma(t\pm vx^{3}),\gamma(vt\pm x^{3})e^{-\frac{1}{2}\eta\gamma^{5}})\ ,$ (83)
@@ 10.2 Jet energy loss and entropy production
Figure 11: Heat–map type plots of the entropy produced by a jet (coming in from the top left in each plot) scattering in a slice of a medium (centre of the figure) for different values of the total charge of the jet in the Schwinger model of Eq. (84) Reprinted[154] under the terms of the Creative Commons 4.0 License, https://creativecommons.org/licenses/by/4.0/
@@ 10.3 State preparation and energy loss in charged medium
Research on state preparation concerns the construction of appropriate ansätze and the minimisation of objective functions such as the expectation value of the energy or the infidelity with respect to a desired target state. Theoretical and practical inputs about the process are based on the preparation and evolution within simplified models, such as the 1+1 D Schwinger model[79], in which the “hadrons” are $e^{+}-e^{-}$ bound states, exemplified with the use of the $112$ qubits quantum-computer ibm_torino[95]. The vacuum of the theory is prepared first, and subsequently a “hadron” on top of it, keeping track of the evolution of the chiral condensate expectation value as the system evolves in time for up to $14$ Trotter steps. The state-preparations procedure is named SC-ADAPT-VQE and consists on repeatedly tuning selected unitary transformation with classical circuits for increasing lattice lengths until convergence, finally extrapolating these transformations to the real length on the quantum device. It is found that classical simulators can manage the preparation for quite large lengths as the correlations between physical sites remain small, but run into trouble for time evolution.
@@
The preparation and time-evolution circuits of[79] are combined with the presence of static charges to mimic, via classical simulators, the evolution of heavy particles in dense media [80]. Lattice effects are probed through the evolution of an initial bound state, where the light dynamical degrees of freedom couple to a heavy charge. Due to the lattice dispersion relations, once a critical velocity $v^{*}$, less than the speed of light, is exceeded, the light degrees of freedom decouple from the charge and excite the lattice in a process resembling hadronisation. To suppress such effects, the velocity of the moving charge is kept small as it traverses the lattice populated with other static charges. Within this setup, the energy loss and entanglement are computed, allowing lattice discretisation effects to be isolated. Entanglement is found to be more sensitive to lattice artifacts than other classical observables. In addition, quantum circuits and associated computational costs are discussed.
10.4 Real Time Scattering
Another application where there is interest is in tracking real–time scattering among hadrons. This is not possible in the lattice, a Euclidean formulation, so it has been thought of [164] as a possible application for quantum computers.
We are a bit less enthusiastic than other authors about this real-time scattering as a flagship application in hadron physics. Indeed, lattice gauge theory is already extracting scattering phase shifts and inelasticities, pole positions, etc. through analysis of the Lüscher formula (via the relation of the density of states at a given energy for fixed angular momentum and the phase shift of the scattering, as lattice methods are very competent at identifying spectral energies). We see the interest of real-time scattering to be more intense in lower-energy fields, such as biochemistry, where there is hope to experimentally follow individual reactants in time. For hadron physics, and for the time being, this is more of a curiosity. This is of course opinionable, so we leave here a mention of the effort and refer the reader to the literature.
@@ 11 Outlook
There are dedicated reviews on Quantum Computing which cover topics in High-Energy Physics [72] and Nuclear Physics [156], but we feel that there is a need to point out the important niche applications in Hadron Physics, straddling both fields, sharing from the large scales of the first but also from the nonperturbative problems of the second. We have striven to give, in a few brushstrokes, a colourful picture of some interesting problems which can be addressed when quantum computers become a more practical tool, without precluding that several other interesting applications may be devised in the field.
@@
This entails that several extant works fall back onto topical phenomenological models addressing one or another aspect of the full theory. As hardware advances, the advantage of directly working with discretisations and truncations of QCD lies in the possibility of controlling systematic uncertainties through well-defined computational parameters—a feature that also applies to quantum simulations.
Finally, one crucial aspect that quantum computing practitioners entering the field of quantum field theory simulations should bear in mind is the renormalisation, or flow, of the Hamiltonian parameters when the scale of the problem changes (and hence the discretisation), whether in momentum space or particle-number space. The following paragraph is devoted to this issue.
Renormalisation Group evolution
We have not dedicated space in this brief appraisal to the renormalisation of the QCD Hamiltonian, but a comment is in order here, since, unlike the quantum mechanical theories often transcribed for a quantum computer, Chromodynamics will require serious work to handle its renormalisation intricacies. A promising line of research is that of the Renormalisation Group Procedure for Effective Particles (RGPEP) [165] recently deployed for a Yukawa Hamiltonian [161]. The traditional idea of continuously transforming the Hamiltonian via a spectrum-conserving unitary rotation parametrised by a control variable $s$ (for example, a scale)
$H(s)=U(s)H(0)U^{\dagger}(s)\ ,$ (85)
@@
New counterterms are then added to the Hamiltonian and the eigenvalues are extracted. This, or other RGE techniques [157, 158, 159] will need to be automated into the workflow for QCD on quantum computers, which is currently best described as simply cutoff (regulated).
The RGPEP is most often formulated using the Light Front, and the Hamiltonian requires renormalisation. The infinite number of degrees of freedom requires truncation; to study the sensitivity thereto, calls for a renormalisation-group method to be applied. The RGPEP enables the construction of effective Hamiltonians through a similarity transformation that depends on a scale parameter. The eigenvector of the effective Hamiltonian contains a small number of non-negligible Fock components, thereby reducing the complexity of the description. Such an approach appears promising for reducing the number of degrees of freedom without truncating terms that encode essential dynamics. Simulations with effective Hamiltonians have been performed for a Yukawa theory within RGPEP, showing that the computational cost to block encode the renormalised Hamiltonian is comparable to block encoding the bare Hamiltonian [160, 161].